# Toward regular black holes in sixth-derivative gravity

### v1

@@ III.3 Solutions with $R=0$ and with $g_{tt}g_{rr}=-1$
To close this section, it is important to mention two classes of solutions that, although prominent in general relativity and quadratic gravity, they do not seem to be very relevant for a generic sixth-derivative gravity model.
@@
Firstly, in fourth-derivative gravity, metrics with vanishing Ricci scalar ($R=0$) played an important role in the identification of certain classes of asymptotically flat solutions [26]. Here, nevertheless, the trivial flat spacetime is the only solution in the form of Eqs. (2) and (3) that satisfies this condition. This can be verified by comparing the solution (LABEL:Sol00Geral) with the expansion of the equation $R=0$ order by order. The lower order equations generate constraints between the free parameters of (LABEL:Sol00Geral) and, at order $r^{8}$, it forces $a_{i}=0=b_{i}$ for $i\geqslant 1$.
@@
Also, solutions such that $g_{tt}g_{rr}=-1$i.e., defined by a single metric function $A(r)=1/B(r)$ — can be shown to have no free parameters, with all the coefficients in (LABEL:Sol00Geral) being determined by the constants in the action. Simply put, we have verified that the constraints between the free parameters in (LABEL:Sol00Geral) are such that $a_{i}=0$ for $i=3,\ldots,10$, but it might happen that $a_{2}$ is a nonzero constant. Although the assumption $A(r)=1/B(r)$ is often employed in the construction of phenomenological models of regular black holes, our result implies that it might be too restrictive in gravity models with more than four derivatives. Indeed, the lack of free parameters makes it difficult to associate solutions of this type with matter sources. This observation might serve as a motivation for further studies of regular effective geometries with $g_{tt}g_{rr}\neq-1$ (see, e.g.[46, 22]).
IV Effect of the terms required by renormalizability
@@
Up to this point the discussion concerned the most general (complete) sixth-derivative gravity model. There are various branches, though, that can be analyzed if some of the coefficients in the action (1) are switched off or taken in particular combinations. A more detailed consideration of these specific scenarios will be carried out in a separate work [41], but here we would like to address the question of the effect of the terms required by renormalizability. In fact, while the action (1) emerges in its most general form from quantum corrections to general relativity, it need not contain all those six-derivative structures if (1) is taken to be the action of a super-renormalizable gravity.
@@
From the point of view of renormalizability, a sixth-derivative gravity model must have sixth-order derivatives in its spin-2 and spin-0 sectors. In terms of the action (1), this corresponds to the requirements $\gamma_{1}\neq 0$ and $3\gamma_{1}+\gamma_{2}\neq 0$. Also, since the counterterms can have up to four metric derivatives [5], for multiplicative renormalizability we must have $\alpha,\beta_{1},\beta_{2}\neq 0$, and a cosmological constant. We omit the latter, for our interest is on the higher derivative’s effects. Therefore, the first five terms in Eq. (1) with the restrictions mentioned above suffice to yield a super-renormalizable model; omitting or including cubic-curvature structures will not affect the renormalizability.
@@
In this spirit, super-renormalizable models can be formulated in terms of other structures quadratic in curvature and with two covariant derivatives, such as $R_{\mu\nu\alpha\beta}\Box R^{\mu\nu\alpha\beta}$ and $C_{\mu\nu\alpha\beta}\Box C^{\mu\nu\alpha\beta}$, because they also yield sixth-order derivatives of the metric. Although they only differ from combinations of $R\Box R$ and $R_{\mu\nu}\Box R^{\mu\nu}$ by cubic and boundary terms, such models are not equivalent if the action does not contain all the terms present in (1).
@@
In Table 1 we summarize the families of solutions that exist around $r=0$ for a super-renormalizable action with all the second- and fourth-derivative structures, and different combinations of sixth-derivative terms (with arbitrary coefficients). Here we restrict considerations to integer values of $s$ and $t$.
Sixth-derivative $(s,t)_{0}$ solution Number of free
@@
$R_{\mu\nu}\Box R^{\mu\nu}$, $C_{\mu\nu\alpha\beta}\Box C^{\mu\nu\alpha\beta}$ $(0,0)_{0}$ $6\to 5$
Table 1: Summary of solutions found around $r=0$. The arrow in the counting of the number of parameters indicates the reduction of parameters after taking into account the freedom to re-scale the time coordinate.
The fourth-derivative gravity also admits solutions with indicial structures $(0,0)_{0}$, $(1,-1)_{0}$ and $(2,2)_{0}$, but the number of free parameters are different [23, 26]. The solutions $(0,0)_{0}$ can be obtained as a particular case of (LABEL:Sol00Geral) by adjusting the coefficients $\gamma_{i}$.
@@
On the other hand, the family $(2,2)_{0}$ here is larger than in fourth-derivative gravity, for it has two more free parameters. These solutions are also singular, with the Kretschmann scalar behaving like $r^{-8}$ as $r\to 0$. However, they are only present in some models constructed with the Weyl tensor. For instance, the Einstein–Hilbert action supplemented by the Goroff–Sagnotti term ${C^{\mu\nu}}_{\alpha\beta}{C^{\alpha\beta}}_{\rho\sigma}{C^{\rho\sigma}}_{\mu\nu}$ also seems to admit a solution $(2,2)_{0}$, characterized by 3 physical free parameters, which we obtained by solving the field equations up to nine orders in $r$.
V Conclusions
@@
There is an important difference between fourth- and sixth-derivative gravity in what concerns the space of static spherically symmetric solutions. In fact, while the former admits families of singular solutions with indicial structure $(1,-1)_{0}$ and $(2,2)_{0}$ [23], in a generic sixth-derivative gravity described by the action (1) (with arbitrary and unrelated couplings) all the solutions of the form of Eqs. (2) and (3) are regular at $r=0$. A similar contrast between these higher-derivative gravities has already been noticed for the linear version of the models [19], and it is rewarding to see that this result has a counterpart at nonlinear level.
@@
We have also identified solutions that contain a horizon. The fact that we only found regular solutions around $r=0$ might suggest that these black hole solutions are regular. In order to confirm this statement, however, it would be necessary to use numerical methods or to prove that the solutions with horizon also have a representation in terms of Frobenius series around the origin. Indeed, there may still exist singular solutions of non-Frobenius type, or the regular solutions at $r=0$ may possess singularities for finite values of $r\neq 0$. We expect that this work will motivate further research on this important topic for the understanding of the role of higher derivatives in gravity.
@@
Regarding incomplete sixth-derivative models, we discussed the influence of the terms required by renormalizability on the space of solutions. In particular, we identified renormalizable models that have singular solutions of the type $(1,-1)_{0}$, and others admitting solutions $(2,2)_{0}$. Nevertheless, since these models are obtained by a special tuning between the coefficients in the general action (1), we conjecture that the solution which is equivalent to Schwarzschild in general relativity is in the class $(0,0)_{0}$ — because it is the one common to all models with sixth-order field equations. The situation here is similar to fourth-derivative gravity, for which the asymptotically flat solution that couples to normal ghost-free matter is in the $(2,2)_{0}$ class, although the model admits Schwarzschild as solution [26].
@@
Last but not least, we found that solutions satisfying the constraints $R=0$ or $g_{tt}g_{rr}=-1$ do not seem to be of much relevance in gravity theories with more than four metric derivatives. Therefore, like in the works [46, 22], it might be necessary to relax those conditions to construct phenomenological models of regular black holes aimed at reproducing aspects of solutions of higher-derivative gravity.
Acknowledgements

### v2

@@ IV Solutions with $R=0$ and with $g_{tt}g_{rr}=-1$
There are two classes of solutions that, although prominent in general relativity and quadratic gravity, do not seem to be very relevant for a generic sixth-derivative gravity model.
@@
In fourth-derivative gravity, metrics with vanishing Ricci scalar ($R=0$) played an important role in the identification of certain classes of asymptotically flat solutions [31]. Here, nevertheless, the trivial flat spacetime is the only solution in the form of Eqs. (3) and (4) that satisfies this condition. This can be verified by comparing the solution (LABEL:Sol00Geral) with the expansion of the equation $R=0$ order by order. The lower-order equations generate constraints between the free parameters of (LABEL:Sol00Geral) and, at order $r^{8}$, it forces $a_{i}=0=b_{i}$ for $i\geqslant 1$.
@@
Also, solutions such that $g_{tt}g_{rr}=-1$ — i.e., defined by a single metric function, $A(r)=1/B(r)$ — can be shown to have no free parameters, with all the coefficients in (LABEL:Sol00Geral) being determined by the constants in the action. Simply put, under these conditions the constraints between the free parameters in (LABEL:Sol00Geral) are such that $b_{i}=0$ for $i=3,4,\ldots$, but it might happen that $b_{2}$ is a nonzero fixed constant. Although the assumption $A(r)=1/B(r)$ is often employed in the construction of phenomenological models of regular black holes, our result implies that it might be too restrictive in gravity models with more than four derivatives, in line with statements in [46, 58]. Indeed, the lack of free parameters makes it difficult to associate solutions of this type with matter sources. This observation might serve as a motivation for further studies of regular geometries with $g_{tt}g_{rr}\neq-1$ (see, e.g., [59, 26]).
V Effect of the terms required by renormalizability
@@
Up to this point, the discussion concerned the most general (complete) sixth-derivative gravity model. There are various branches, though, that can be analyzed if some of the coefficients in the action (1) are switched off or taken in particular combinations. A more detailed consideration of these specific scenarios will be carried out in a separate work [45], but here we would like to address the question of the effect of the terms required by renormalizability. In fact, while the action (1) emerges in its most general form from quantum corrections to general relativity, it need not contain all those six-derivative structures if it is taken to be the action of a superrenormalizable gravity.
@@
From the point of view of renormalizability, a sixth-derivative gravity model must have sixth-order derivatives in its spin-2 and spin-0 sectors. In terms of the action (1), this corresponds to the requirements $\gamma_{2}\neq 0$ and $3\gamma_{1}+\gamma_{2}\neq 0$. Also, since the counterterms can have up to four metric derivatives [7], for multiplicative renormalizability we must have $\alpha,\beta_{1},\beta_{2}\neq 0$, and a cosmological constant. We omit the latter, for our interest is in the higher derivative’s effects. Therefore, the terms in the first line of Eq. (1) with the restrictions mentioned above suffice to yield a superrenormalizable model; omitting or including cubic-curvature structures will not affect the renormalizability.
@@
In this spirit, superrenormalizable models can be formulated in terms of other structures quadratic in curvature and with two covariant derivatives, such as $R_{\mu\nu\alpha\beta}\Box R^{\mu\nu\alpha\beta}$ or $C_{\mu\nu\alpha\beta}\Box C^{\mu\nu\alpha\beta}$, because they also yield sixth-order derivatives of the metric. Although they only differ from combinations of $R\Box R$ and $R_{\mu\nu}\Box R^{\mu\nu}$ by cubic and boundary terms, such models are not equivalent if the action does not contain all the terms present in (1).
@@
In Table 1 we summarize the families of solutions that exist around $r=0$ for a superrenormalizable action with all the two- and four-derivative structures and different combinations of six-derivative terms (with arbitrary coefficients). Here we restrict considerations to integer values of $s$ and $t$.
Six-derivative $(s,t)_{0}$ solution Number of free
@@
Table 1: Summary of solutions around $r=0$. The arrow in the counting of the number of parameters indicates the reduction of parameters after taking into account the freedom to rescale the time coordinate.
Some models admit a family of solutions with indicial structure $(1,-1)_{0}$ or $(2,2)_{0}$, in addition to the family $(0,0)_{0}$. The latter is described as a particular case of (LABEL:Sol00Geral), by adjusting the coefficients $\gamma_{i}$ according to the terms present in the action. Solutions with indicial structures $(0,0)_{0}$, $(1,-1)_{0}$, and $(2,2)_{0}$ also occur in the fourth-derivative gravity, but with a different number of free parameters [27, 31].
@@
On the other hand, the family $(2,2)_{0}$ here is larger than in fourth-derivative gravity, for it has two more free parameters. These solutions are also singular, with the Kretschmann scalar behaving like $r^{-8}$ as $r\to 0$. However, they are only present in some models constructed with the Weyl tensor [60].
VI Conclusions
@@
There is an important difference between fourth- and sixth-derivative gravity in what concerns the space of static spherically symmetric solutions. In fact, while the former admits families of singular solutions, with indicial structure $(1,-1)_{0}$ and $(2,2)_{0}$ [27], in a generic sixth-derivative gravity described by the action (1) (with arbitrary and unrelated couplings) all the solutions of the form of Eqs. (3) and (4) belong to the family $(0,0)_{0}$ and are regular at $r=0$. This possibility was suggested in [28], and our analysis corroborates the result. A similar contrast between these higher-derivative gravities had already been noticed for the linear version of the models [21], and it is rewarding to see that it has a counterpart at the nonlinear level.
@@
We have also identified solutions that contain a horizon. The fact that we only found regular solutions around $r=0$ might suggest that these black hole solutions are regular. In order to confirm this statement, however, it would be necessary to use numerical methods or to prove that the solutions with a horizon also have a representation in terms of Frobenius series around the origin. Indeed, there may still exist singular solutions of non-Frobenius type, or regular solutions at $r=0$ may possess singularities for finite values of $r\neq 0$. The study of numerical solutions to sixth-derivative gravity is still in the early stages (see, e.g., [61]), and we expect that our work will motivate further research on this important topic for the understanding of the role of higher derivatives in gravity.
@@
Regarding incomplete sixth-derivative models, we discussed the influence of the terms required by renormalizability on the space of solutions. In particular, we identified renormalizable models that have singular solutions of the type $(1,-1)_{0}$, and others admitting solutions $(2,2)_{0}$. Nevertheless, since these models are obtained by a special tuning between the coefficients in the general action (1), we conjecture that the solution that is equivalent to the Schwarzschild one in general relativity is in the class $(0,0)_{0}$ — because it is the one common to all such models with sixth-order field equations. The situation here is similar to fourth-derivative gravity, for which the asymptotically flat solution that couples to normal ghost-free matter is in the $(2,2)_{0}$ class, although the model admits the Schwarzschild black hole as solution [31].
@@
Last but not least, we found that solutions satisfying the constraints $R=0$ or $g_{tt}g_{rr}=-1$ do not seem to be of much relevance in gravity theories with more than four metric derivatives (with generic couplings). Therefore, like in [26], it might be necessary to relax those conditions to construct phenomenological models of regular black holes aimed at reproducing aspects of solutions of higher-derivative gravity.
Acknowledgements.
@@
Appendix A System of indicial equations and its solution
The system of indicial equations for the Frobenius series solutions of sixth-derivative gravity around $r=0$ is obtained by substituting the ansatz for the metric, Eqs. (3) and (4), into the field equations originated from the action (1) and expanding to the lowest order in $r$. The resulting expressions have the general form of Eq. (6), but here we shall split the analysis into three cases: $s>0$, $s=0$, and $s<0$.
A.1 Case $s>0$
If $s>0$, the expansion of the field equations at lowest order yields
$\frac{H_{tt}(r)}{B(r)}$ $=$ $\frac{1}{32a_{0}^{3}\,r^{6+3s}}\sum_{i=1}^{8}\gamma_{i}\,g_{i}(s,t)+O\left(\frac{1}{r^{5+3s}}\right),$
$H_{rr}(r)$ $=$ $\frac{1}{32a_{0}^{2}\,r^{6+2s}}\sum_{i=1}^{8}\gamma_{i}\,h_{i}(s,t)+O\left(\frac{1}{r^{5+2s}}\right),$ (16)
where the coefficients $g_{i}(s,t)$ and $h_{i}(s,t)$ depend on the quantities $s$ and $t$ through
$g_{1}(s,t)$ $=$ $4(s+2)[s(t+4)-t(t+2)-4]\big{[}(6s-t)\big{(}10s^{2}-st-t^{2}\big{)}+2\big{(}118s^{2}-34st-t^{2}\big{)}+4(69s-14t+22)\big{]},$ (17)
$h_{1}(s,t)$ $=$ $4(s+2)[s(t+4)-t(t+2)-4]\big{[}t\big{(}10s^{2}-st-t^{2}\big{)}+2\big{(}20s^{2}+6st-3t^{2}\big{)}+4(31s-6t+14)\big{]},$ (18)
$g_{2}(s,t)$ $=$ $2\big{[}st\big{(}60s^{4}-76s^{3}t+11s^{2}t^{2}+6st^{3}-t^{4}\big{)}+2\big{(}60s^{5}+146s^{4}t-208s^{3}t^{2}+35s^{2}t^{3}+8st^{4}-t^{5}+328s^{4}+152s^{3}t$ (19)
$-263s^{2}t^{2}+96st^{3}-3t^{4}\big{)}+8\big{(}124s^{3}-69s^{2}t-15st^{2}+21t^{3}-5s^{2}-110st+2t^{2}\big{)}-96(9s+2t+3)\big{]},$
$h_{2}(s,t)$ $=$ $2\big{[}st^{2}\big{(}10s^{3}-11s^{2}t+t^{3}\big{)}+2t\big{(}20s^{4}+6s^{3}t-21s^{2}t^{2}+2st^{3}+t^{4}\big{)}+2\big{(}60s^{4}+56s^{3}t-71s^{2}t^{2}-4st^{3}+21t^{4}\big{)}$ (20)
$+8\big{(}64s^{3}-9s^{2}t-33st^{2}-7t^{3}+47s^{2}-26st+6t^{2}\big{)}+96(2t-5s-3)\big{]},$
$g_{3}(s,t)$ $=$ $4[s(t+4)-t(t+2)-4]^{2}\big{[}30s^{2}+st-t^{2}+2(47s-t+37)\big{]},$ (21)
$h_{3}(s,t)$ $=$ $4[s(t+4)-t(t+2)-4]^{2}[t(5s+t)+2(10s+4t+23)],$ (22)
$g_{4}(s,t)$ $=$ $2\big{[}t^{2}(s-t)^{2}\big{(}30s^{2}+st-t^{2}\big{)}+2t\big{(}80s^{4}-55s^{3}t-71s^{2}t^{2}+47st^{3}-t^{4}\big{)}+2\big{(}140s^{4}+146s^{3}t-193s^{2}t^{2}+22st^{3}$ (23)
$+25t^{4}\big{)}+8\big{(}59s^{3}-20s^{2}t+23st^{2}+12t^{3}-39s^{2}+28st+56t^{2}\big{)}-32(8s-21t-19)\big{]},$
$h_{4}(s,t)$ $=$ $2\big{[}t^{3}(s-t)^{2}(5s+t)+8t^{2}\big{(}5s^{3}-6s^{2}t+t^{3}\big{)}+2t\big{(}70s^{3}-35s^{2}t-34st^{2}+27t^{3}\big{)}+8\big{(}30s^{3}+5s^{2}t-23st^{2}+14t^{3}$ (24)
$+19s^{2}-28st+48t^{2}\big{)}-32(21s-16t-17)\big{]},$
$g_{5}(s,t)$ $=$ $t^{2}(s-t)^{2}\big{(}30s^{2}+st-t^{2}\big{)}+12st\big{(}10s^{3}-3s^{2}t-14st^{2}+7t^{3}\big{)}+6\big{(}40s^{4}+84s^{3}t-23s^{2}t^{2}-8st^{3}+7t^{4}\big{)}$ (25)
$+8\big{(}71s^{3}+69s^{2}t+24st^{2}+8t^{3}\big{)}+24\big{(}6s^{2}+16st+11t^{2}\big{)}-32(3s-9t-10),$
$h_{5}(s,t)$ $=$ $t^{3}(s-t)^{2}(5s+t)+6t^{2}\big{(}5s^{3}-5s^{2}t-st^{2}+t^{3}\big{)}+6t\big{(}20s^{3}+3s^{2}t-12st^{2}+9t^{3}\big{)}+8\big{(}25s^{3}+27s^{2}t-3st^{2}$ (26)
$-5t^{3}\big{)}+24\big{(}10s^{2}+8st+9t^{2}\big{)}-32(9s-15t-14),$
$g_{6}(s,t)$ $=$ $t^{2}(s-t)^{2}\big{(}30s^{2}+st-t^{2}\big{)}+2t\big{(}40s^{4}+19s^{3}t-97s^{2}t^{2}+37st^{3}+t^{4}\big{)}+4\big{(}20s^{4}+24s^{3}t-22s^{2}t^{2}-43st^{3}$ (27)
$+11t^{4}\big{)}+8\big{(}10s^{3}-5s^{2}t+23st^{2}-10t^{3}-25s^{2}+54st+39t^{2}\big{)}-32(5s-18t-7),$
$h_{6}(s,t)$ $=$ $t^{3}(s-t)^{2}(5s+t)+4t^{2}\big{(}5s^{3}-3s^{2}t-3st^{2}+t^{3}\big{)}+4st\big{(}10s^{2}-9st-3t^{2}\big{)}+8\big{(}20s^{3}-7s^{2}t-14st^{2}+19t^{3}$ (28)
$+11s^{2}-14st+21t^{2}\big{)}-32(15s-4t-11),$
$g_{7}(s,t)$ $=$ $4\big{[}t^{2}(s-t)^{2}\big{(}30s^{2}+st-t^{2}\big{)}+2t\big{(}40s^{4}+19s^{3}t-97s^{2}t^{2}+37st^{3}+t^{4}\big{)}+2\big{(}40s^{4}+128s^{3}t+43s^{2}t^{2}-92st^{3}$ (29)
$+21t^{4}\big{)}+16\big{(}15s^{3}-3s^{2}t+27st^{2}-4t^{3}\big{)}+8\big{(}46s^{2}-36st+45t^{2}\big{)}+32(15s+13)\big{]},$
$h_{7}(s,t)$ $=$ $4\big{[}t^{3}(s-t)^{2}(5s+t)+4t^{2}\big{(}5s^{3}-3s^{2}t-3st^{2}+t^{3}\big{)}+2t\big{(}20s^{3}+17s^{2}t-20st^{2}+11t^{3}\big{)}+8\big{(}20s^{3}-2s^{2}t$ (30)
$+5st^{2}-t^{3}+26s^{2}-8st+31t^{2}\big{)}-32(6s-3t-11)\big{]},$
$g_{8}(s,t)$ $=$ $4\big{[}t^{2}(s-t)^{2}\big{(}30s^{2}+st-t^{2}\big{)}+6t^{2}\big{(}31s^{3}-41s^{2}t+9st^{2}+t^{3}+70s^{2}-54st+4t^{2}\big{)}-8\big{(}s^{3}-66st^{2}+19t^{3}\big{)}$ (31)
$+32\big{(}9t^{2}+1\big{)}\big{]},$
$h_{8}(s,t)$ $=$ $4\big{[}t^{3}(s-t)^{2}(5s+t)+24st^{3}(s-t)+12t^{3}(3s-t)+8\big{(}10s^{3}-3s^{2}t+7t^{3}\big{)}+144s^{2}+160\big{]}.$ (32)
Hence, for the field equations to be solved at lowest order in $r$ for any values of $\gamma_{1,\ldots,8}$, we must have
$g_{i}(s,t)=h_{i}(s,t)=0,\quad\forall\,i=1,\ldots,8,$ (33)
which constitute the system of indicial equations for $s>0$.
To prove that the system (33) does not admit solutions with $s>0$, consider the following observations:
  • i.
    The subsystem $g_{3}(s,t)=h_{3}(s,t)=0$ only admits the one-parameter family of solutions
    $s=\frac{t^{2}+2t+4}{t+4},\quad t\neq-4.$ (34)
    Indeed, it is straightforward to verify that
    $h_{3}(s,t)=0\quad\Longleftrightarrow\quad s=\frac{t^{2}+2t+4}{t+4}\quad\text{or}\quad s=-\frac{t^{2}+8t+46}{5(t+4)}.$ (35)
    The latter option yields
    $g_{3}(s(t),t)=-\frac{1728(t-2)(t^{2}+3t+11)^{3}}{25(t+4)^{2}}=0,\quad t\in\mathbb{R}\quad\Longleftrightarrow\quad t=2,$ (36)
    forcing $s=-11/5<0$, which violates the assumption on the sign of $s$. On the other hand, the former option, given by Eq. (34), solves $g_{3}(s,t)=0$ for any value of $t\neq-4$.
  • ii.
    The only admissible solution for the subsystem $g_{3}(s,t)=h_{3}(s,t)=0$ and $h_{4}(s,t)=0$ is $s=-t=1$.
    In fact, substituting (34) into the expression for $h_{4}(s,t)$ we obtain
    $h_{4}(s,t)=\frac{64(t+1)^{2}(t^{2}+3t+11)(t^{2}+4t+12)}{(t+4)^{2}},$ (37)
    whose roots can be easily obtained using Bhaskara’s formula. The only real solution is $t=-1$, which yields $s=1$.
Finally, it is immediate to verify that $s=-t=1$ is not a zero of any of the functions in Eqs. (29)–(32). This completes the proof that the system (33) does not have solutions with $s>0$.
A.2 Case $s=0$
In this case the expansions of $H_{tt}$ and $H_{rr}$ have the form
$\frac{H_{tt}(r)}{B(r)}$ $=$ $\frac{1}{32a_{0}^{3}\,r^{6}}\sum_{i=1}^{8}\gamma_{i}\,g_{i}(0,t,a_{0})+O\left(\frac{1}{r^{5}}\right),$
$H_{rr}(r)$ $=$ $\frac{1}{32a_{0}^{2}\,r^{6}}\sum_{i=1}^{8}\gamma_{i}\,h_{i}(0,t,a_{0})+O\left(\frac{1}{r^{5}}\right),$ (38)
now with coefficients $g_{i}(0,t,a_{0})$ and $h_{i}(0,t,a_{0})$ that also depend on $a_{0}$, namely,
$g_{1}(0,t,a_{0})$ $=$ $-8\big{[}4(1-a_{0})+t(t+2)\big{]}\big{[}t^{3}-2t^{2}-56t+8(a_{0}+11)\big{]},$ (39)
$h_{1}(0,t,a_{0})$ $=$ $8\big{[}4(1-a_{0})+t(t+2)\big{]}\big{[}t^{3}+6t^{2}+24t-8(a_{0}+7)\big{]},$ (40)
$g_{2}(0,t,a_{0})$ $=$ $-4\big{[}t^{5}+3t^{4}-84t^{3}-8(a_{0}+1)t^{2}+96(a_{0}+1)t-48\big{(}a_{0}^{2}+2a_{0}-3\big{)}\big{]},$ (41)
$h_{2}(0,t,a_{0})$ $=$ $4\big{[}t^{5}+21t^{4}-28t^{3}+8(a_{0}+3)t^{2}+32(3-5a_{0})t+48\big{(}a_{0}^{2}+2a_{0}-3\big{)}\big{]},$ (42)
$g_{3}(0,t,a_{0})$ $=$ $-4\big{[}4(1-a_{0})+t(t+2)\big{]}^{2}\big{[}t^{2}+2t+2(a_{0}-37)\big{]},$ (43)
$h_{3}(0,t,a_{0})$ $=$ $4\big{[}4(1-a_{0})+t(t+2)\big{]}^{2}\big{[}t^{2}+8t+2(a_{0}+23)\big{]},$ (44)
$g_{4}(0,t,a_{0})$ $=$ $-2\big{[}t^{6}+2t^{5}-2(a_{0}+25)t^{4}-8(a_{0}+12)t^{3}+32(5a_{0}-14)t^{2}+672(a_{0}-1)t+32(a_{0}-1)^{2}(a_{0}-19)\big{]},$ (45)
$h_{4}(0,t,a_{0})$ $=$ $2\big{[}t^{6}+8t^{5}-2(a_{0}-27)t^{4}+8(14-3a_{0})t^{3}+32(12-7a_{0})t^{2}+32\big{(}a_{0}^{2}-17a_{0}+16\big{)}t+32(a_{0}-1)^{2}(a_{0}+17)\big{]},\,\,\,\,\,\,\,\,\,\,\,\,$ (46)
$g_{5}(0,t,a_{0})$ $=$ $-\big{[}t^{6}-42t^{4}-64t^{3}-24(a_{0}+11)t^{2}+288(a_{0}-1)t+32(a_{0}-1)^{2}(a_{0}-10)\big{]},$ (47)
$h_{5}(0,t,a_{0})$ $=$ $t^{6}+6t^{5}+54t^{4}-40t^{3}-24(a_{0}-9)t^{2}-480(a_{0}-1)t+32(a_{0}-1)^{2}(a_{0}+14),$ (48)
$g_{6}(0,t,a_{0})$ $=$ $-\big{[}t^{6}-2t^{5}-44t^{4}-8(a_{0}-10)t^{3}+24(a_{0}-13)t^{2}+576(a_{0}-1)t+32(a_{0}-1)^{2}(a_{0}-7)\big{]},$ (49)
$h_{6}(0,t,a_{0})$ $=$ $t^{6}+4t^{5}-8(a_{0}-19)t^{3}-168(a_{0}-1)t^{2}-128(a_{0}-1)t+32(a_{0}-1)^{2}(a_{0}+11)$ (50)
$g_{7}(0,t,a_{0})$ $=$ $-4\big{[}t^{6}-2t^{5}-2(a_{0}+21)t^{4}+8(a_{0}+8)t^{3}+72(a_{0}-5)t^{2}+32(a_{0}-1)^{2}(a_{0}-13)\big{]},$ (51)
$h_{7}(0,t,a_{0})$ $=$ $4\big{[}t^{6}+4t^{5}-2(a_{0}-11)t^{4}-8(a_{0}+1)t^{3}+8(31-7a_{0})t^{2}+32\big{(}a_{0}^{2}-4a_{0}+3\big{)}t+32(a_{0}-1)^{2}(a_{0}+11)\big{]},$ (52)
$g_{8}(0,t,a_{0})$ $=$ $-4\big{[}t^{6}-6t^{5}-24t^{4}+152t^{3}-288t^{2}+32(a_{0}-1)^{3}\big{]},$ (53)
$h_{8}(0,t,a_{0})$ $=$ $4\big{[}t^{6}-12t^{4}+56t^{3}+32(a_{0}-1)^{2}(a_{0}+5)\big{]}.$ (54)
Therefore, for the field equations to be solved at lowest order in $r$ for any values of $\gamma_{1,\ldots,8}$, we must have
$g_{i}(0,t,a_{0})=h_{i}(0,t,a_{0})=0,\quad\forall\,i=1,\ldots,8.$ (55)
Notice that this system of indicial equations for $s=0$ might also act as constraint to the value of the parameter $a_{0}$ in Eq. (3).
To prove that the only solution of (55) is $t=0$ and $a_{0}=1$, we proceed like in the previous case and start by considering the coefficients related to the parameter $\gamma_{3}$. We have the following results:
  • i.
    The only admissible solutions of the subsystem $g_{3}(0,t,a_{0})=h_{3}(0,t,a_{0})=0$ are either $t=-20$ and $a_{0}=-143$ or in the form
    $a_{0}=\frac{1}{4}\left(t^{2}+2t+4\right),\quad t\in\mathbb{R}.$ (56)
    In fact, it is straightforward to verify that
    $h_{3}(0,t,a_{0})=0\quad\Longleftrightarrow\quad a_{0}=\frac{1}{4}\left(t^{2}+2t+4\right)\quad\text{or}\quad a_{0}=-\frac{1}{2}\left(t^{2}+8t+46\right).$ (57)
    The latter option yields
    $g_{3}(0,t,a_{0}(t))=216(t+20)(t^{2}+6t+32)^{2}=0,\quad t\in\mathbb{R}\quad\Longleftrightarrow\quad t=-20,$ (58)
    which results in $a_{0}=-143$. On the other hand, the former solution, Eq. (56), solves $g_{3}(0,t,a_{0})=0$ for any real $t$.
  • ii.
    The only admissible solution for the subsystem $g_{3}(0,t,a_{0})=h_{3}(0,t,a_{0})=0$ and $h_{4}(0,t,a_{0})=0$ is $t=0$ and $a_{0}=1$.
    Indeed, by direct substitution one can verify that $t=-20$ and $a_{0}=-143$ does not solve $h_{4}(0,t,a_{0})=0$. As for solutions in the form (56), substituting into (46) we obtain
    $h_{4}(0,t,a_{0}(t))=2t^{2}\big{(}t^{2}+2\big{)}(t^{2}+6t+32),$ (59)
    whose only real root is $t=0$, which yields $a_{0}=1$.
It is immediate to verify that $t=0$ and $a_{0}=1$ is a zero of all the functions (39)–(54), being, therefore, the only solution of the system (55).
A.3 Case $s<0$
Finally, the lowest-order term of the expansion of $H_{tt}$ and $H_{rr}$ for negative values of $s$ reads
$\frac{H_{tt}(r)}{B(r)}$ $=$ $-\frac{4\gamma_{3}+2\gamma_{4}+\gamma_{5}+\gamma_{6}+4\gamma_{7}+4\gamma_{8}}{r^{6}}+O\big{(}r^{-5}\big{)},$
$H_{rr}(r)$ $=$ $\frac{4\gamma_{3}+2\gamma_{4}+\gamma_{5}+\gamma_{6}+4\gamma_{7}+4\gamma_{8}}{r^{6-s}}a_{0}+O\big{(}r^{s-5},r^{-6}\big{)}.$
Therefore, the field equations cannot be solved at lowest order for arbitrary values of the parameters $\gamma_{i}$ if $s<0$.